OpenStudy (anonymous):

Solve. 3x2 + 12x = 3

OpenStudy (anonymous):

i bet you can do this divide by 3 first

OpenStudy (anonymous):

then, because the middle term is going to have an even coefficient (namely 4) solve by completing the square

OpenStudy (anonymous):

divide 12 by three?

OpenStudy (anonymous):

i can walk you through it if you like or else you can try first

OpenStudy (anonymous):

I'll come up with an answer first :)

OpenStudy (anonymous):

divide everything on both sides of the equal sign by 3, everything

OpenStudy (anonymous):

The expression is the same as: \[3x^2+12x−3=0\] Now notice they all can divide by 3, we get \[x^2+4x−1=0\] Now use the technique I taught you in your last 'finding zero' question

OpenStudy (anonymous):

Nevermind, use the quadratic formula

OpenStudy (anonymous):

-2 _+_ sq root 3

OpenStudy (anonymous):

nop :(

OpenStudy (anonymous):

\[x^2+4x=1\]\[(x+2)^2=1+4=5\] \[x+2=\pm\sqrt{5}\]

OpenStudy (anonymous):

ughh let me try again!!!1

OpenStudy (anonymous):